In one day, Sweatsmartz, inc. sold 125 tee-shirts. The blue v-neck ones sold for $7.95 each and the red crew neck ones sold for $8.95. The company took in $1036.75. How many of each kind were sold?
Answer by stanbon(75887)

In one day, Sweatsmartz, inc. sold 125 tee-shirts. The blue v-neck ones sold for $7.95 each and the red crew neck ones sold for $8.95. The company took in $1036.75. How many of each kind were sold?
--------------
Quantity:: b + r = 125 shirts
Value:::795b+895r = 103675 cents
----------------------------
Modify for elimination:
795b + 795r = 795*125
795b + 895r = 103675
--------------------------
Subtract and solve for "r":
100r = 4300
r = 43 (# of red sold)
-----
b = 125-43 = 82 (# of blue sold)
----------
Cheers,
Stan H.
============